OpenStudy (anonymous):

f(x)=-2x^2+2x-3, what is f(5), please explain

OpenStudy (anonymous):

We just need to substitute 5 for every x in the given function f(x), and simplify to get the value of f(5).

OpenStudy (anonymous):

\[f(x)=-2x^2+2x-3\] \[f(5)=-2(5)^2+2(5)-3\] \[f(5)=-2(25)+10-3\] \[f(-5)=-50+10-3\] \[f(-5)=-43\]

OpenStudy (anonymous):

Thanks so much

OpenStudy (anonymous):

You're welcome. Thank you.